# The stochastic discrete nonlinear Schrödinger equation: microscopic derivation and finite-temperature phase transition

### v1

@@ Appendix E The harmonic chain
$Z_{\beta}=\pi^{L}$ $\int_{\mathbb{R}_{+}^{L-1}}\exp\left(-\beta\sum_{\nu=1}^{L-1}\omega^{2}_{q_{\nu}}\left|\hat{c}_{q_{\nu}}\right|^{2}\right)$ (85)
$\Theta\left(1-\sum_{\nu=1}^{L-1}\left|\hat{c}_{q_{\nu}}\right|^{2}\right)\prod_{\nu=1}^{L-1}d\left|\hat{c}_{q_{\nu}}\right|^{2}$
@@
$\zeta(|\beta|)=$ $\int_{\mathbb{R}_{+}^{L-1}}\exp\left(-\eta\sum_{\nu=1}^{L-1}\omega_{q_{\nu}}^{2}y_{\nu}\right)$ (87)
$\Theta\left(|\beta|-\sum_{\nu=1}^{L-1}y_{\nu}\right)\prod_{\nu=1}^{L-1}dy_{\nu}\,.$
@@
$\frac{1}{L}\sum_{\nu=0}^{L-1}\ln(s+\eta 2(1-\cos(q_{\nu})))$
$\simeq$ $\frac{1}{2\pi}\int_{0}^{2\pi}\ln(s+\eta 2(1-\cos(q)))dq$
@@
$\mathcal{L}(nI_{n}(t)/t)(s)$ $=$ $\left(s-\sqrt{s^{2}-1}\right)^{n}$
$=$ $\frac{1}{\left(s+\sqrt{s^{2}-1}\right)^{n}},\quad\mbox{Re}(s)>1$
@@
$E(\beta)$ $=$ $-\frac{\partial\ln Z_{\beta}}{\partial\beta}=2-2\eta\frac{I_{L+1}(2|\beta|)}{I_{L}(2|\beta|)}$ (99)
$=$ $2-2\eta\frac{I_{L+1}(L\cdot 2|\beta|/L)}{I_{L}(L\cdot 2|\beta|/L)}$

### v2

@@ Appendix E The harmonic chain
$Z_{\beta}=\pi^{L}\int_{\mathbb{R}_{+}^{L-1}}\exp\left(-\beta\sum_{\nu=1}^{L-1}\omega^{2}_{q_{\nu}}\left|\hat{c}_{q_{\nu}}\right|^{2}\right)\Theta\left(1-\sum_{\nu=1}^{L-1}\left|\hat{c}_{q_{\nu}}\right|^{2}\right)\prod_{\nu=1}^{L-1}d\left|\hat{c}_{q_{\nu}}\right|^{2}$ (84)
@@
$\zeta(|\beta|)=\int_{\mathbb{R}_{+}^{L-1}}\exp\left(-\eta\sum_{\nu=1}^{L-1}\omega_{q_{\nu}}^{2}y_{\nu}\right)\Theta\left(|\beta|-\sum_{\nu=1}^{L-1}y_{\nu}\right)\prod_{\nu=1}^{L-1}dy_{\nu}\,.$ (86)
@@
$\frac{1}{L}\sum_{\nu=0}^{L-1}\ln(s+\eta 2(1-\cos(q_{\nu})))$ $\simeq$ $\frac{1}{2\pi}\int_{0}^{2\pi}\ln(s+\eta 2(1-\cos(q)))dq$
@@
$\mathcal{L}(nI_{n}(t)/t)(w)$ $=$ $\left(w-\sqrt{w^{2}-1}\right)^{n}=\frac{1}{\left(w+\sqrt{w^{2}-1}\right)^{n}},\quad\mbox{Re}(w)>1$ (93)
@@
$E(\beta)=-\frac{\partial\ln Z_{\beta}}{\partial\beta}=2-2\eta\frac{I_{L+1}(2|\beta|)}{I_{L}(2|\beta|)}=2-2\eta\frac{I_{L+1}(L\cdot 2|\beta|/L)}{I_{L}(L\cdot 2|\beta|/L)}$ (98)